OpenStudy (anonymous):

help i will fan and medal A coin is tossed and a number cube is rolled. what is p(heads, a number less that 5)?

OpenStudy (anonymous):

@xGuardians

OpenStudy (anonymous):

\[A. \frac{ 1 }{ 3 }\] \[B. \frac{ 5 }{ 12 }\] \[C. \frac{ 2 }{ 3 }\] \[D. \frac{ 5 }{ 6 }\]

OpenStudy (jim766):

1/2 * 4/6 = 4/12 = 1/3

OpenStudy (anonymous):

My best guess is 1/3

OpenStudy (alexwee123):

there's 0.5 chance of getting a heads on the coin and 4 numbers that are less than 5 on a die so we'd get 1/2 * 4/6= 1/2 * 2/3 = 1/3

OpenStudy (anonymous):

ok thats guys